Sunrise, Inc., has no debt outstanding and a total market value of $332,100. Earnings before interest and taxes, EBIT, are projected to be $48,000 if economic conditions are normal. If there is strong expansion in the economy, then EBIT will be 18 percent higher. If there is a recession, then EBIT will be 29 percent lower. The company is considering a $170,000 debt issue with an interest rate of 7 percent. The proceeds will be used to repurchase shares of stock. There are currently 8,100 shares outstanding. The company has a tax rate of 21 percent, a market-to-book ratio of 1.0, and the stock price remains constant.

  

a-1.

Calculate earnings per share (EPS) under each of the three economic scenarios before any debt is issued. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

a-2. Calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
b-1. Calculate earnings per share (EPS) under each of the three economic scenarios assuming the company goes through with recapitalization. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
b-2. Given the recapitalization, calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
a-1. Recession EPS
Normal EPS
Expansion EPS
a-2. Recession percentage change in EPS -29.00 %
Expansion percentage change in EPS 18.00 %
b-1. Recession EPS
Normal EPS
b-2. Expansion EPS
Recession percentage change in EPS -38.56 %
Expansion percentage change in EPS 23.93 %
Answer #1
a-1
EPS = EBIT*(1-tax rate)/shares outstanding
Recession
EPS = EBIT*(1-recession impact%)*(1-tax rate)/shares outstanding
EPS=48000*(1-0.29)*(1-0.21)/8100
EPS=3.32
Normal
EPS = EBIT*(1-tax rate)/shares outstanding
EPS=48000*(1-0.21)/8100
EPS=4.68
Expansion
EPS = EBIT*(1+Growth impact%)*(1-tax rate)/shares outstanding
EPS=48000*(1+0.18)*(1-0.21)/8100
EPS=5.52
a-2
%age change in EPS for Recession
=(EPS recession/EPS normal-1)*100
=(3.3239/4.6815-1)*100
=-29%
%age change in EPS for Growth
=(EPS Growth/EPS normal-1)*100
=(5.5241/4.6815-1)*100
=18%
b-1
New no. of shares = old shares-debt/(Market value/old shares)
=8100-170000/(332100/8100)
=3954
EPS = (EBIT-debt*interest%)*(1-tax rate)/new shares outstanding
Recession
EPS = (EBIT*(1-recession impact%)-debt*interest %age)*(1-tax rate)/new shares outstanding
EPS=(48000*(1-0.29)-170000*0.07)*(1-0.21)/3954
EPS=4.43
Normal
EPS = (EBIT-debt*interest%)*(1-tax rate)/new shares outstanding
EPS=(48000-170000*0.07)*(1-0.21)/3954
EPS=7.21
Expansion
EPS = (EBIT*(1+growth impact%)-debt*interest %age)*(1-tax rate)/new shares outstanding
EPS=(48000*(1+0.18)-170000*0.07)*(1-0.21)/3954
EPS=8.94
b-2
%age change in EPS for Recession
=(EPS recession/EPS normal-1)*100
=(4.4315/7.2127-1)*100
=-39%
%age change in EPS for Growth
=(EPS Growth/EPS normal-1)*100
=(8.9389/7.2127-1)*100
=24%